# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v1

@@ Abstract
We show that the three commonly employed approaches that define the same intrinsic linear anomalous Hall response actually lead to different results for intrinsic nonlinear transport. The difference arises from an intrinsic anomalous distribution. It originates from scattering, but its value is completely independent of scattering, because it represents the local equilibration of electron wave packets with field corrected energy. As a manifestation, we find that under ac driving, the intrinsic contributions in rectified component and in double-frequency component exhibit distinct frequency dependence, which can be probed in experiment. Using first-principles calculations, we estimate the signals that can be probed in antiferromagnetic CuMnAs.
@@
Intrinsic responses, which depend only on band structures of materials, are of crucial importance in condensed matter physics, since they can be evaluated accurately with first-principles techniques, offer benchmarks for experiments, and probe interesting geometric band properties. A prominent example is anomalous Hall effect [1, 2], whose intrinsic contribution manifests the Berry curvature of band structures [3, 4]. In that context, there have been three different approaches to define what is the intrinsic response in diffusive transport. (1) The first approach is to consider the so-called “clean limit”, without any scattering. For example, in the semiclassical theory [5, 2], this means one directly takes the equilibrium distribution $f_{0}$ in the calculation of response. (2) The second is to define the intrinsic response as the extrapolation of the ac Hall response from the high-frequency regime ($\omega\tau\gg 1$, where $\tau$ is the scattering time) to the dc limit, with $1/\tau\rightarrow 0$ faster than $\omega\rightarrow 0$ [1]. (3) The third one is an operational and intuitive definition, namely, one fully considers scattering and defines intrinsic contribution as the scattering independent component of the total response. For anomalous Hall transport, the three approaches are considered equivalent, in the sense that they yield the same intrinsic response.
Recently, there has been tremendous interest in nonlinear Hall effect, for which the charge current response $j\sim E^{2}$ is quadratic in the applied $E$ field [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Under broken time reversal symmetry, this effect also allows an intrinsic contribution [6, 29, 40]. Despite intensive studies, the nature of this intrinsic nonlinear transport has not been thoroughly understood. Particularly, it remains unclear whether the three approaches above still define the same intrinsic response in the nonlinear regime. If not, which one will be a more proper definition?
In this work, we address the above fundamental problem. We find that distinct from linear response, the three approaches actually lead to three different results for second order nonlinear transport (see Table. 1). This difference originates from the field correction to the electron wave packet energy, which must be included for nonlinear response as it is quadratic in $E$. Owing to this energy shift, in the presence of disorder scattering, electrons acquire a correction in distribution function, the intrinsic anomalous distribution, which turns out to be completely independent of scattering. Physically, the corrected distribution can be naturally interpreted as the equilibrium state for an ensemble of electron wave packets with the corrected energy. The establishment of this new equilibrium requires scattering, but the resulting distribution does not depend on scattering. It follows that approach 1 must differ from approach 3, since it does not consider any scattering effect; meanwhile, approach 2 partially captures the above contribution and leads to a result sitting in-between. We further investigate the situation under ac driving. Interestingly, we find that even within the intraband transport domain, the intrinsic rectified response $j^{(0)}$ and the intrinsic double-frequency response $j^{(2\omega)}$ have different amplitudes in approach 3. In addition, $j^{(2\omega)}$ also exhibits different values in low-frequency and high-frequency regimes. Based on these insights, we argue that approach 3 fits into a natural and consistent physical picture and also connects closely with experimental measurement, so it appears to be a more appropriate definition. We also discuss possible ways to test some of our predictions by nonlinear measurement on antiferromagnetic CuMnAs.
Table 1: Different results of second order intrinsic nonlinear transport under the three definitions of intrinsic contribution which are equivalent for linear response.
@@
Component without counting scattering $\ \omega\tau\gg 1$ and $\omega\rightarrow 0$  with scattering, expressions independent of disorder
dc limit    Eq. (6)  $\frac{1}{2}$(Eq. (6)+Eq. (7))    Eq. (7)
@@
ac double-frequency    Eq. (6)   --    $\omega\tau\gg 1$, Eq. (6);  $\omega\tau\ll 1$, Eq. (7)
Preliminaries. In the semiclassical theory, the charge current density is expressed as (we set $e=\hbar=1$)
@@
where $l\equiv(n,\bm{k})$ is a composition of band and crystal momentum labels, and $\bm{v}_{l}$ is the velocity of the electron wave packet. Up to second order in applied uniform $E$ field, the scattering-independent part of velocity is given by [6, 42]
@@
is the field-corrected wave packet energy: Besides the unperturbed band energy $\varepsilon_{l}$, there is, as mentioned, a second order correction [43, 44, 37]
@@
where $V^{nn^{\prime}}$ is the interband velocity matrix element, and the quantities on the right hand side are evaluated at the same $k$ point. In Eq. (2), the $k$-space Berry curvature $\tilde{\bm{\Omega}}_{l}$ includes the first-order field correction, i.e., $\tilde{\bm{\Omega}}=\bm{\Omega}+\delta\bm{\Omega}$, where $\bm{\Omega}$ is for unperturbed bands, and interestingly, the correction $\delta\Omega_{a}=\epsilon_{abc}\partial_{k_{b}}G_{cd}E_{d}$ ($\epsilon_{abc}$ is the totally antisymmetric tensor) is also determined by BCP [6, 23, 24].
To obtain the intrinsic nonlinear current, in approach 1, one simply put $f_{l}=f_{0}(\varepsilon_{l})$, substitute Eq. (2) into (1), and collect terms of $E^{2}$ order. The result is
@@
One may already spot an ambiguity in the treatment above, namely, since we are effectively dealing with a bunch of field-dressed wave packets with energies $\tilde{\varepsilon}_{l}$, it seems more appealing to use their equilibrium distribution $f_{0}(\tilde{\varepsilon}_{l})$ in Eq. (1) for the intrinsic current. By doing so and performing the expansion $f_{0}(\tilde{\varepsilon})=f_{0}({\varepsilon})+f_{0}^{\prime}({\varepsilon})\delta\varepsilon$, we obtain a different and rather simple intrinsic nonlinear current:
@@
which is a purely Hall (hence dissipationless) response. The two currents differ by
@@
is the difference between the two distributions, which we refer to as the intrinsic anomalous distribution.
In fact, $\bm{j}^{\mathrm{in}}$ in Eq. (7) has been obtained in several pioneering works on this effect [6, 23, 24]. We shall uncover the scattering origin of its difference from $\bm{j}^{cl}$ and see below that it in fact corresponds to the intrinsic response defined by approach 3.
The role of scattering. Now, the question is: Where does the intrinsic anomalous distribution $g_{l}$ come from? The original system in the absence of driving field is in the equilibrium state described by $f_{0}(\varepsilon_{l})$. When one gradually turns on the $E$ field, the electron wave packets are dressed, with their energies changed to $\tilde{\varepsilon}_{l}$. Clearly, $f_{0}(\varepsilon_{l})$ makes no longer an equilibrium state for particles with these renormalized energies, so it should relax towards an (new) equilibrium corresponding to $f_{0}(\tilde{\varepsilon}_{l})$. This is the part of the distribution function in addition to the non-equilibrium part due to the drift of Fermi surface by field and other extrinsic effects. Evidently, the relaxation from $f_{0}(\varepsilon_{l})$ to $f_{0}(\tilde{\varepsilon}_{l})$ must be through scattering.
The picture above can be justified by examining the Boltzmann equation, which describes the kinetics of electron wave packets. Assuming uniform system and field, it reads $\partial_{t}f_{l}-\bm{E}\cdot\partial_{\bm{k}}{f_{l}}=\mathcal{I}\{f_{l}\}$, where the right hand side is the collision integral, acting as a linear operator on $f_{l}$. Note that the $E$-field driving term here produces the non-equilibrium part $\sim E\tau$ in the distribution function, which does not contribute to the intrinsic current. Therefore, to study intrinsic response, we may drop this term and focus on the solution of
@@
Note that this equation is describing wave packets with energies $\tilde{\varepsilon}_{l}$. In other words, the scattering processes in $\mathcal{I}$ are defined for these dressed wave pacekts. For example, in the case of one-electron elastic scattering, we have [45]
@@
with the scattering rate $\omega_{l^{\prime}l}=2\pi|T_{l^{\prime}l}|^{2}\delta(\tilde{\varepsilon}_{l^{\prime}}-\tilde{\varepsilon}_{l}),$ where $T$ is the scattering $T$-matrix [2]. Then, it is clear that $f_{0}(\tilde{\varepsilon}_{l})$ is always a static solution of (10) (an explicit treatment of this simple case can be found in the Supplemental Material [46]). In contrast, $f_{0}({\varepsilon}_{l})$ is not a static solution: If imposed as an initial distribution, it should evolve by Eq. (10) and relax through scattering towards the static solution $f_{0}(\tilde{\varepsilon}_{l})$ (In this process, one may assume that chemical potential and temperature of the system are fixed by environment).
The above analysis is general, regardless of the details of the system as well as scattering. It follows that there always exists the intrinsic anomalous distribution $g_{l}$ in the dc limit, which originates from scattering but its value is independent of scattering. Thus, in approach 3, the intrinsic nonlinear current should be evaluated with $f_{0}(\tilde{\varepsilon}_{l})$ instead of $f_{0}({\varepsilon}_{l})$, and the result indeed takes the form of Eq. (7).
Frequency dependence. It remains to find the intrinsic response defined by approach 2. In this approach, we need to consider the frequency dependence due to ac driving. Let us take $\tilde{E}_{a}(t)=E_{a}\cos\omega t=\text{Re}(E_{a}e^{i\omega t})$. At $E^{2}$ order, the current response contains two frequency components: $j_{a}=\text{Re}(j_{a}^{(0)}+j_{a}^{(2\omega)}e^{i2\omega t})$, a zero-frequency (rectified) component $j^{(0)}$ and a double-frequency component $j^{(2\omega)}$. Both components may contain intrinsic contributions.
@@
It is known that the $g^{(\omega)}$ component contributes only to extrinsic terms [20], so we focus on $g^{(0)}$ and $g^{(2\omega)}$.
From the analysis of Eq. (10) just now, one can readily see that the zero-frequency component $g^{(0)}$ always contains the following intrinsic anomalous distribution
@@
where $\delta\varepsilon^{(0)}_{l}$ denotes the zero-frequency component of the energy correction. This intrinsic anomalous distribution is independent of frequency. Here, we assumed $\omega<\Delta$ with $\Delta$ the size of local gap at Fermi surface, such that the intraband transport dominates the response.
Meanwhile, for the double-frequency component $g^{(2\omega)}$, its value depends on frequency and differs in two regimes. In the low-frequency regime $\omega\ll\tau^{-1}$, the left hand side of (10) is negligible, so the result takes the same form as the zero-frequency component:
@@
This can be readily understood, since at such high frequency, the oscillating electrons simply do not “see” the scatterers. For frequency between the two regimes, $g_{l}^{(2\omega)}$ should exhibit explicit $\omega$ dependence, and crossover from one regime to the other. In the Supplemental Material [46], we show that for the case of elastic scattering with relaxation time approximation for the Boltzmann equation, the crossover takes the form of
@@
Inputting these results into Eq. (1), one finds that, in the sense of approach 3, for the low-frequency regime $\omega\ll\tau^{-1}$, the ac intrinsic nonlinear current is given by
@@
where $j^{\mathrm{in}}_{a}$ is given by expression (7) with $E$ the amplitude of ac $E$ field. As for the high-frequency regime $\tau^{-1}\ll\omega<\Delta$, the result is
@@
Therefore, the double-frequency intrinsic nonlinear response exhibits difference between low and high frequency regimes, whereas the rectified response does not. This frequency dependence is a unique feature of nonlinear response that has not been revealed before. It has important experimental implications that will be discussed below.
@@
Now, if we use approach 2 to define the dc intrinsic nonlinear current, i.e., taking Eq. (18) and extrapolating to dc limit, the result is given by $(1/2)(\bm{j}^{\mathrm{cl}}+\bm{j}^{\mathrm{in}})$. Thus, the three approaches give three different results for the intrinsic nonlinear response. All these results are summarized in Table 1.
@@
Discussion. We have revealed that the three approaches, which define the same intrinsic linear anomalous Hall response, produce different results when applied to the nonlinear effect. The question is which one is more reasonable. Their difference comes from the intrinsic anomalous distribution function $g_{l}$. As we have shown, although $g_{l}$ has its root in scattering, its value is completely independent of scattering. It arises from the natural equilibration of electrons under a local shift of energy. It follows that in dc measurement, one can never separate $g_{l}$ from $f_{0}(\varepsilon_{l})$; in other words, they always come together in the form of $f_{0}(\tilde{\varepsilon}_{l})$ in producing the response. Thus, for better comparison between theory and experiment, approach 3 seems to be the more proper definition.
It is worth noting that the nature of this intrinsic anomalous distribution contribution is different from the so-called side jump contribution in anomalous Hall effect [47, 2]. In early literature, it is often stated that the side jump contribution is independent of scattering. However, this conclusion is not valid in general: it has been shown that the side jump contribution depends on the type of disorders and generally varies in the presence of more than one scattering sources [48, 49]. In contrast, the intrinsic anomalous distribution is completely independent of these details, so it can be regarded as an intrinsic response.
@@
The result from approach 3, such as Eq. (7), is also physically appealing. In the picture of semiclassical theory, we are dealing with dynamics of a bunch of electron wave packets. The external fields not only supply driving force, they also dress the wave packets, leading to their renormalized energy $\tilde{\varepsilon}_{l}$ and Berry curvature $\tilde{\Omega}$. Then, in a consistent treatment, it is natural that the response is formulated in terms of these dressed wave packets with their renormalized properties. In this sense, the use of $f_{0}(\tilde{\varepsilon}_{l})$ for intrinsic response is a physically consistent choice. In the same spirit, for a consistent treatment, the scattering amplitude $|T_{l^{\prime}l}|^{2}$ should be defined with respect to the dressed wave packets, which has been shown in Ref. [50].
The same logic applies to the case under applied magnetic field. It is well known that to linear order in $B$, the wave packet energy acquires a correction $\delta\varepsilon_{l}=-\bm{m}_{l}\cdot\bm{B}$, where $\bm{m}_{l}$ is the spin plus orbital magnetic moment of the wave packet [51, 5]. Then, the study of magneto-responses should also include the intrinsic anomalous distribution due to this correction $\delta\varepsilon_{l}$. Notably, it has been shown that failing to account for this intrinsic anomalous distribution will lead to the unphysical result that a static $B$ field alone can drive a transport current [52, 53, 54]. This is a strong evidence showing the necessity to include the intrinsic anomalous distribution for a consistent theory.

### v2

@@ Abstract
We show that the three commonly employed approaches that define the same dc (or low-frequency) intrinsic linear anomalous Hall response actually lead to different results for intrinsic nonlinear transport. The difference is due to an intrinsic anomalous distribution (IAD). It originates from a nonlinear field effect during scattering, but its value is completely independent of scattering, because it represents the local equilibration of electron wave packets with field corrected energy. The proper definition of intrinsic current that is detectable in experiment must incorporate the effect of IAD. We also show that IAD is indispensable for consistency with fundamental physical relations. In addition, we predict that under ac driving, the intrinsic reponses in rectified and double-frequency channels exhibit distinct frequency dependence, for which we estimate the signals that can be probed in antiferromagnetic CuMnAs.
@@
Intrinsic responses, which depend only on band structures of materials, are of crucial importance in condensed matter physics, since they can be evaluated accurately with first-principles techniques, offer benchmarks for experiments, and probe interesting geometric band properties. A prominent example is anomalous Hall effect [1, 2], whose intrinsic contribution manifests the Berry curvature of band structures [3, 4]. In that context, there have been three different approaches to define what is intrinsic response in dc diffusive transport. (D1) The first approach is to consider the naive “clean limit”, without any scattering. For example, in the semiclassical framework [5, 2], this means one directly takes the equilibrium distribution $f_{0}$ in the calculation of response. (D2) The second is to define intrinsic response as the extrapolation of the ac Hall response from the high-frequency regime ($\omega\gg\tau^{-1}$, where $\tau$ is the scattering time) to the dc limit, with $\tau^{-1}\rightarrow 0$ faster than $\omega\rightarrow 0$ [1]. (D3) The third one is an operational definition, namely, one fully considers scattering and defines intrinsic response as the scattering independent component of the total response. For linear anomalous Hall transport, the three approaches yield the same result hence are considered equivalent.
Recently, there has been tremendous interest in nonlinear Hall effect, for which the charge current response $j\propto E^{2}$ is quadratic in the driving $E$ field [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Under broken time reversal symmetry, this effect also allows an intrinsic response [6]. Despite intensive studies, there remains controversy on the very definition of this intrinsic response, which has caused conflicting interpretations of experimental results. For example, two pioneering experiments Ref. [30] and Ref. [31], both on even-layer MnBi2Te4, are interpreted with two differently defined intrinsic nonlinear currents [corresponding to (D3) and (D1), respectively, as we see below] which even possess distinct symmetry characters. This issue of proper definition of intrinsic nonlinear current has generated lasting confusion and become a central problem of nonlinear transport, which requires an urgent solution.
Table 1: Different results of intrinsic nonlinear current obtained from the three definitions. The second, third, and fourth columns correspond to (D1), (D2), and (D3), respectively.
@@
Component without counting scattering $\ \omega\tau\gg 1$ and $\omega\rightarrow 0$  with scattering, expressions independent of scattering
dc response    Eq. (6)  $\frac{1}{2}$(Eq. (6)+Eq. (7))    Eq. (7)
@@
ac double-frequency    Eq. (6)   --    $\omega\ll 1/\tau$, Eq. (7);  $\omega\gg 1/\tau$, Eq. (6)
In this work, we address this critical problem. We find that distinct from linear transport, the three approaches actually lead to different results for nonlinear transport (see Table 1). This difference can be traced to the field correction to the electron wave packet energy, which is quadratic in $E$. Owing to this energy shift, in the presence of disorder scattering, electrons acquire a correction in distribution function, termed as the intrinsic anomalous distribution (IAD), which turns out to be completely independent of scattering. Physically, the corrected distribution which includes IAD can be naturally interpreted as the equilibrium state for an ensemble of electron wave packets with the corrected energy. It is crucial that the establishment of this new equilibrium requires scattering, but the resulting distribution itself has no dependence on scattering at all. It follows that for dc transport, (D3) fully captures IAD and gives a purely Hall response; (D1) must differ from (D3), since it does not consider any scattering effect; meanwhile, (D2) partially captures IAD and leads to a result sitting in-between. We argue that (D3) is the proper definition of intrinsic nonlinear response, because IAD is an inseparable part of intrinsic response that is actually detected in dc or low-frequency experiment. In other words, (D3) is the appropriate one that should be used for interpreting experiments.
Furthermore, the notion of IAD is universal, and its significance is not limited to nonlinear transport. We demonstrate that it is indispensable for any coherent theory which maintains consistency with fundamental physics relations and principles, such as the Einstein relation and the absence of dc current driven by a static uniform magnetic field.
In addition, as a by-product of this new understanding, we predict that under ac driving, the second-harmonic component of intrinsic nonlinear current may acquire different values in low- and high-frequency regimes. Using first-principles calculation, we estimate signals that can be detected in CuMnAs.
Preliminaries. Consider diffusive transport under a dc driving $E$ field. The charge current density is (we set $e=\hbar=1$)
@@
where $l\equiv(n,\bm{k})$ is a composition of band and crystal momentum labels, and $\bm{v}_{l}$ is the velocity of electron wave packet. Up to second order in $E$, the scattering-independent part of velocity is given by [6, 44]
@@
is the field-corrected wave packet energy: Besides the unperturbed band energy $\varepsilon_{l}$, there is a second-order field correction [45, 46, 40]
@@
where $V^{nn^{\prime}}$ is the interband velocity matrix element, and quantities on both sides are evaluated at the same $k$ point. In Eq. (2), the Berry curvature $\tilde{\bm{\Omega}}_{l}$ includes the first-order field correction, i.e., $\tilde{\bm{\Omega}}=\bm{\Omega}+\delta\bm{\Omega}$, where $\bm{\Omega}$ is for unperturbed bands, and the correction $\delta\Omega_{a}=\epsilon_{abc}\partial_{k_{b}}G_{cd}E_{d}$ ($\epsilon_{abc}$ is the totally antisymmetric tensor) is also determined by BCP [6, 24, 25].
To obtain the intrinsic nonlinear current, in approach (D1), one simply put $f_{l}=f_{0}(\varepsilon_{l})$, substitute Eq. (2) into (1), and collect terms of $E^{2}$ order. The result is
@@
One may already spot an ambiguity in the treatment above, namely, since we are effectively dealing with a bunch of field-dressed wave packets with energies $\tilde{\varepsilon}_{l}$, it seems more appealing to use their equilibrium distribution $f_{0}(\tilde{\varepsilon}_{l})$ in Eq. (1) for the intrinsic current. By doing so and performing the expansion $f_{0}(\tilde{\varepsilon})=f_{0}({\varepsilon})+f_{0}^{\prime}({\varepsilon})\delta\varepsilon$, one obtains a different and simpler intrinsic current [47]:
@@
which is a purely Hall (hence dissipationless) response.
The two currents differ by
@@
is the difference between the two distributions, which we refer to as IAD.
$\bm{j}^{\mathrm{in}}$ in Eq. (7) has been obtained in several pioneering works on this effect [6, 24, 25]. We shall show that it corresponds to approach (D3) in dc and low-frequency transport. IAD is the key for the difference between $\bm{j}^{\mathrm{in}}$ [from (D3)] and $\bm{j}^{\mathrm{cl}}$ [from (D1)]. It must be emphasized that this difference is dramatic: The two results even have distinct symmetry characters, e.g., $\bm{j}^{\mathrm{in}}$ is purely Hall and is forbidden by any rotational symmetry normal to the transport plane (such as $C_{3z}$) [30], whereas $\bm{j}^{\mathrm{cl}}$ is not.
Intrinsic anomalous distribution. Now, the question is: Where does the IAD $g_{l}$ come from?
The original system in the absence of driving field is in the equilibrium state described by $f_{0}(\varepsilon_{l})$. When one gradually turns on the $E$ field, the electron wave packets are dressed, with their energies changed to $\tilde{\varepsilon}_{l}$. Clearly, $f_{0}(\varepsilon_{l})$ makes no longer an equilibrium state for particles with these renormalized energies, so it should relax towards an (new) equilibrium corresponding to $f_{0}(\tilde{\varepsilon}_{l})$. This is the part of the distribution function in addition to the non-equilibrium part due to the drift of Fermi surface by field and other extrinsic effects. Evidently, the relaxation from $f_{0}(\varepsilon_{l})$ to $f_{0}(\tilde{\varepsilon}_{l})$ must be through scattering.
The physical picture above can be justified by examining the Boltzmann equation, which describes the kinetics of electron wave packets. Assuming uniform system and field, it reads $\partial_{t}f_{l}-\bm{E}\cdot\partial_{\bm{k}}{f_{l}}=\mathcal{I}\{f_{l}\}$, where the right hand side is the collision integral, acting as a linear operator on $f_{l}$. The $E$-field driving term here produces the usual non-equilibrium part $\propto E\tau$ in the distribution function, which, however, does not contribute to the intrinsic response. Therefore, to study intrinsic response, one may drop this term and focus on the solution of
@@
Note that this equation is describing wave packets with energies $\tilde{\varepsilon}_{l}$. In other words, the scattering processes in $\mathcal{I}$ are defined for these dressed wave pacekts. For example, for impurity scattering, we may write [48]
@@
with the scattering rate $\omega_{l^{\prime}l}=2\pi|T_{l^{\prime}l}|^{2}\delta(\tilde{\varepsilon}_{l^{\prime}}-\tilde{\varepsilon}_{l}),$ where $T$ is the scattering $T$-matrix [2]. Then, it is clear that $f_{0}(\tilde{\varepsilon}_{l})$ is always a static solution of (10) (an explicit treatment of this simple case can be found in [47]). In contrast, $f_{0}({\varepsilon}_{l})$ is not a static solution: If imposed as an initial distribution, it should evolve by Eq. (10) and relax through scattering towards the static solution $f_{0}(\tilde{\varepsilon}_{l})$. In this process, their difference, i.e., IAD, comes about.
The above analysis is general, regardless of the details of the system as well as scattering. It follows that there always exists the IAD $g_{l}$ in dc transport, which originates from scattering but its value is completely independent of scattering. Thus, in (D3), the intrinsic nonlinear current should be evaluated with $f_{0}(\tilde{\varepsilon}_{l})$ instead of $f_{0}({\varepsilon}_{l})$, and the result indeed takes the form of Eq. (7).
Frequency dependence. It remains to find the result from (D2). In this approach, we need to consider the frequency dependence under ac driving. Take $\tilde{E}_{a}(t)=E_{a}\cos\omega t=\text{Re}(E_{a}e^{i\omega t})$. At $E^{2}$ order, the current response contains two components:
$j_{a}=\text{Re}(j_{a}^{(0)}+j_{a}^{(2\omega)}e^{i2\omega t}),$ (12)
a zero-frequency (rectified) component $j^{(0)}$ and a second-harmonic component $j^{(2\omega)}$. Both may contain intrinsic contributions.
@@
It is known that $g^{(\omega)}$ contributes only to extrinsic terms [20], so we focus on $g^{(0)}$ and $g^{(2\omega)}$.
From the analysis of Eq. (10) just now, one can readily see that the rectified component $g^{(0)}$ always contains the following IAD
@@
where $\delta\varepsilon^{(0)}_{l}$ denotes the zero-frequency component of the energy correction. This IAD is independent of frequency. Here, we assumed $\omega<\Delta$ with $\Delta$ the size of local gap at Fermi surface, such that the intraband transport dominates the response.
Meanwhile, for the second-harmonic component $g^{(2\omega)}$, its value depends on frequency and differs in two regimes. In the low-frequency regime $\omega\ll\tau^{-1}$, the left hand side of (10) is negligible, so the result takes the same form as the zero-frequency component:
@@
This can be readily understood, since at such high frequency, the oscillating electrons simply do not “see” the scatterers. For frequency between the two regimes, $g_{l}^{(2\omega)}$ should exhibit explicit $\omega$ dependence, and crossover from one regime to the other. In [47], we show that under relaxation time approximation, the crossover takes the form of
@@
Inputting these results into Eq. (1), one finds that, in the sense of (D3), for the low-frequency regime $\omega\ll\tau^{-1}$, the ac intrinsic nonlinear current is given by
@@
where $j^{\mathrm{in}}_{a}$ is given by expression (7) with $E$ the amplitude of ac $E$ field. On the other hand, for the high-frequency regime $\tau^{-1}\ll\omega<\Delta$, the result is
@@
Therefore, the second-harmonic response exhibits difference between low and high frequency regimes, whereas the rectified response does not. This is a unique feature of nonlinear response that has not been revealed before. Its experimental implications will be discussed in a while.
Now, we can find the result of dc intrinsic nonlinear current defined by (D2). This is by taking Eq. (19) and extrapolating it to dc limit. The result is given by $(1/2)(\bm{j}^{\mathrm{cl}}+\bm{j}^{\mathrm{in}})$, which, in a sense, only partially captures IAD.
@@
Therefore, the three approaches, which are considered equivalent for linear transport, give three different results for intrinsic nonlinear current. All these results are summarized in Table 1.
@@
Which is the proper definition? We have shown that the three approaches, which define the same intrinsic linear anomalous Hall response, produce distinct results when applied to nonlinear transport. The question is: Which one is appropriate? As clarified, their difference comes from IAD $g_{l}$, which has its root in scattering, but its contribution is completely independent of scattering. It arises from the natural local equilibration of electrons under a local shift of energy. It follows that in dc or low-frequency measurement, one can never separate $g_{l}$ from $f_{0}(\varepsilon_{l})$; they always come together in the form of $f_{0}(\tilde{\varepsilon}_{l})$ in producing the intrinsic response that is detected in experiment. IAD is an inseparable part of this response. This shows (D3) is the proper definition of intrinsic nonlinear current, since it corresponds to what is really measured in experiment.
@@
The result from approach (D3), with $f_{0}(\tilde{\varepsilon}_{l})$ for dc (low-frequency) transport, also has a clear physical picture and is physically appealing. In the picture of semiclassical theory, we are dealing with dynamics of a bunch of electron wave packets. The external fields not only supply driving force, they also dress the wave packets, leading to their renormalized energy $\tilde{\varepsilon}_{l}$ and Berry curvature $\tilde{\Omega}$. Then, in a consistent treatment, it is natural that the response is formulated in terms of these dressed wave packets with their renormalized properties. In this sense, the use of $f_{0}(\tilde{\varepsilon}_{l})$ for intrinsic response is a physically consistent choice. In the same spirit, for a consistent treatment, the scattering amplitude $|T_{l^{\prime}l}|^{2}$ should also be defined with respect to the dressed wave packets, which has been shown in Ref. [49].
Our finding settles the controversy surrounding the experiments [30, 31]. It indicates that the theoretical interpretation adopted in Ref. [30] is correct. The intrinsic nonlinear response probed in the experiment should correspond to Eq. (7), which is purely Hall and forbidden by $C_{3z}$ symmetry, as verified in [30]. Meanwhile, the interpretation [using Eq. (6)] for experiment [31] is incorrect. The nonlinear signal that scales as $\sigma_{xx}^{0}$ found in that experiment should not be attributed to intrinsic response from MnBi2Te4 layers, because of the preserved $C_{3z}$. It is more likely to be attributed to so-called zeroth order extrinsic contributions [50], as supported by several recent studies [50, 51, 52].
Discussion. In this study, we have revealed the significant role of IAD in nonlinear transport. It is worth noting that the nature of IAD is different from the so-called side-jump extrinsic contribution in anomalous Hall effect [2]. In early literature [1], it is often stated that side-jump contribution is independent of scattering. However, this statement is invalid in general: side jump depends on the type of disorders and generally varies in the presence of more than one scattering sources [53, 54]. In contrast, IAD is completely independent of these details, so it should be included in intrinsic response.